xx PunctureTracker

PunctureTracker

Michael Koppitz <koppitz@aei.mpg.de>

Date

Abstract

PunctureTracker tracks black hole Punctures (singularities) by subtracting the value of the shift at the current location (the pucnture is advected with the shift) storing the location in a “SphericalSurface” one can use this information in Regrid2 to let the grid follow the BHs.

1 Parameters

PunctureTracker has the followoing options: -ptxtrackxpunctures (boolean, that triggers action) -ptxverbose (integer that give the verbosity level) -ptxnumxtracked (integer that tells how many Punctures are to be tracked) -ptxinitialxx (double: Initial x coordinate positions of BHs) -ptxinitialxy (double: Initial y coordinate positions of BHs) -ptxinitialxz (double: Initial z coordinate positions of BHs) -ptxwhichxsurfacextoxtake (integer: which Spherical Surface should be used to store the new puncture location)

2 Grid Scalars -(ptxlocxx, ptxlocxy, ptxlocxz) location of puncture -(ptxshiftx, ptxshifty, ptxshiftz) value of shift at location of puncture

3 Parameters

11initialxx 10Scope: privatex01 REAL

11Description: Initial x coordinate positions of punctures

11  Range  10  x01  Default: 0.0

11*:*  10

11

11initialxy  10Scope: privatex01  REAL

11Description: Initial y coordinate positions of punctures

11  Range  10  x01  Default: 0.0

11*:*  10

11

11initialxz  10Scope: privatex01  REAL

11Description: Initial z coordinate positions of punctures

11  Range  10  x01  Default: 0.0

11*:*  10

11

11modifyxdistance  10Scope: privatex01  REAL

11Description: Modify levels when the distance is less than this

11 Range 10 x01 Default: 0.0
110.0:*  10zero or positive

11

11modifyxpuncture  10Scope: privatex01  INT

11Description: Punctures to use for modification criteria

11 Range 10 x01 Default: -1
11-1:9  10One of the tracking punctures or negative for no modification

11

11newxreflevelxnumber 10Scope: privatex01  INT

11Description: The new number of refinement levels

11 Range 10 x01 Default: -1
11-1:*  10Negative for no change

11

11shiftxlimit  10Scope: privatex01  REAL

11Description: Shift components must be less than this in magnitude

11 Range 10 x01 Default: 1.0e10
110.0:*  10zero or positive

11track  10Scope: privatex01  BOOLEAN

11Description: Track this puncture

11                    10  x01  Default: no

11

11verbose  10Scope: privatex01  BOOLEAN

11Description: speak up?

11                    10  x01  Default: no

11

11whichxsurfacextoxstorexinfo 10Scope: privatex01  INT

11Description: A spherical surface index where we can store the puncture location

11 Range 10 x01 Default: -1
11-1  10don’t store puncture location
110:*  10any spherical surface index

11

11nsurfaces  10Scope: shared from SPHERICALSURFACEx01INT

11

4 Interfaces

General Implements: puncturetracker

Inherits: admbase carpetregrid2 sphericalsurface

Grid Variables

4.0.1 PRIVATE GROUPS

11x Group Names x10x Variable Names xx01 Details x11x
11  ptxloc  10  x01  compact110
11                 10  ptxlocxt  x01 description11Location of punctures
11                 10  ptxlocxx  x01 dimensions110
11                 10  ptxlocxy  x01 distribution11CONSTANT
11                 10  ptxlocxz  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1110
11                 10  x01variable type11REAL
11  ptxlocxp  10  x01  compact110
11                 10  ptxlocxtxp  x01 description11Previous location of punctures
11                 10  ptxlocxxxp  x01 dimensions110
11                 10  ptxlocxyxp  x01 distribution11CONSTANT
11                 10  ptxlocxzxp  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1110
11                 10  x01variable type11REAL
11  ptxvel  10  x01  compact110
11                 10  ptxvelxt  x01 description11Velocity of punctures
11                 10  ptxvelxx  x01 dimensions110
11                 10  ptxvelxy  x01 distribution11CONSTANT
11                 10  ptxvelxz  x01 group type11SCALAR
11                 10  x01 timelevels111
11                 10  x01vararrayxsize1110
11                 10  x01variable type11REAL

11

5 Schedule This section lists all the variables which are assigned storage by thorn EinsteinAnalysis/PunctureTracker. Storage can either last for the duration of the run (Always means that if this thorn is activated storage will be assigned, Conditional means that if this thorn is activated storage will be assigned for the duration of the run if some condition is met), or can be turned on for the duration of a schedule function.

Storage x

11Always:  11x
11ptxloc ptxlocxp ptxvel11x

11x  11x

11

Scheduled Functions

CCTKxINITIAL x puncturetrackerxinit x calculate initial location of punctures x

11x11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function

11

CCTKxEVOL x puncturetrackerxtrack x calculate new location of punctures x

11x11After: 11molxevolution
11 x 11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function

11

CCTKxPOSTSTEP x puncturetrackerxsetpositions x copy puncture positions to spherical surfaces x

11x11Language:11c
11 x 11Options: 11global
11 x 11Type: 11function